Fill in the blank: The point where all of the rays will cross each other is known as the refraction X
Answer #1

Focus

Given that, D=30$ cm$; d=0.00039$ mm$; \lambda=820 $ nm$;

Angular Width of central maxima can be given as

d\sin\theta=2\lambda\Rightarrow d\times\frac{w}{D}=2\times820\times10^{-9}\\ \Rightarrow w=\frac{2\times820\times10^{-9}\times30}{0.00039\times10^{-2}}=12.61$ cm$